A rocket burns fuel at a rate of 294 KG/S and exhaust gas at a relative speed of 6KM/S find the thrust of the rocket answer in units of MN
1
Expert's answer
2020-12-02T09:50:38-0500

The thrust of the rocket can be found from the formula:


"T=v\dfrac{dm}{dt},"

here, "T" is the thrust of the rocket, "v=6\ \dfrac{km}{s}" is the relative speed of the exhaust gas, "\dfrac{dm}{dt}=294\ \dfrac{kg}{s}" is the mass flow rate (or the rate of the burn fuel).

Then, we get:


"T=6\ \dfrac{km}{s}\cdot \dfrac{1000\ m}{1\ km}\cdot294\ \dfrac{kg}{s}=1.764\cdot10^6\ N=1.764\ MN."

Answer:

"T=1.764\cdot10^6\ N=1.764\ MN."